OpenStudy (anonymous):

What is the volume of the cone? 12 cm3 20 cm3 36 cm3 60 cm3

OpenStudy (anonymous):

cone

OpenStudy (anonymous):

I don't know the formula

Directrix (directrix):

The volume of a cone is 1/3 Bh where B is the area of the base. A cone has a circular base so its area is π r². That gives V = 1/3 * π r²* h

Directrix (directrix):

V = 1/3 * π r²* h V = 1/3 π 3²* 4 V = ? @Joshuam1 It looks as it you may need to use 3.14 as an approximation for pi. I'll check my answer against yours after you post what you get. Okay?

OpenStudy (anonymous):

okay :D

Directrix (directrix):

Got an answer yet?

OpenStudy (anonymous):

i did 1/3 x b x h

OpenStudy (anonymous):

the base is 3 and the height is 4 right?

Directrix (directrix):

The base radius is 3. The base area is pi*3*3

Directrix (directrix):

You need to do this: 1/3 * π 3²* 4

OpenStudy (anonymous):

oh xD so it's 1/3 x 3.14 x 3 x 3 x 4?

Directrix (directrix):

And, that is what?

OpenStudy (anonymous):

37.68

Directrix (directrix):

That is what I got but it is not one of the options although it is close. Check the problem and the options. Be sure we were not asked to find surface area.

OpenStudy (anonymous):

no, that is all the closest answer is 36

OpenStudy (anonymous):

I have another question but it wants to find the surface area of a cone

Directrix (directrix):

If all that checks out, go with the 36 is all I know to do.

OpenStudy (anonymous):

okay :D

OpenStudy (anonymous):

here is the cone for the surface area question. should i just ask a whole new question or just stay in this one?

Directrix (directrix):

Surface Area of Cone = π*r*L+ π *r^2 where r is the radius of the base and L is the slant height.

OpenStudy (anonymous):

3.14 x 4 x 5 + 3.14 x 4 x 4 ?

Directrix (directrix):

Yes, you are correct. Crank it out.

OpenStudy (anonymous):

113.04

Directrix (directrix):

I used more pi digits and got 113.0973. No big deal.

OpenStudy (anonymous):

oh okay

Directrix (directrix):

Is that an option or are those options still off as they were on the last problem?

OpenStudy (anonymous):

still off 60 cm2 20 cm2 16 cm2 36 cm2

Directrix (directrix):

Off is putting it mildly. Aren't those the volume options again? We're doing surface area now.

OpenStudy (anonymous):

they gave similar options not sure why

Directrix (directrix):

Surface area includes the base. Lateral area does not include the base. Let's see what the lateral area is.

Directrix (directrix):

3.14 x 4 x 5 is the lateral area. What does that equal?

OpenStudy (anonymous):

62.8

OpenStudy (anonymous):

that's a more realistic answer

Directrix (directrix):

You need to talk with the teacher about this. Surface area is 113 ish. Lateral area is 62.8 ish. Both are off the options as was the volume. We should not have to guess what the options are if our work is correct and the problem is correctly stated.

Directrix (directrix):

When you talk with the teacher, have your work handy.

OpenStudy (anonymous):

yes you are right I will have a talk with her about it and see what is up with that

OpenStudy (anonymous):

next question?

OpenStudy (anonymous):

5 more I need help with

Directrix (directrix):

Put it in a new thread. Once a thread gets longer than 10 to 15 rectangles, this OS program gets herky-jerky on my computer. Something about scripts.

OpenStudy (anonymous):

oh alright :)